OpenStudy (anonymous):

Divide zero point 3 2 divided by 8 . 4 forty zero point zero 4 zero point 4

OpenStudy (anonymous):



OpenStudy (anonymous):

wat do u mean @Auxuris

OpenStudy (anonymous):

Multiply eleven point 9 times twenty-two point 1. 2 point 6 2 9 9 two thousand six hundred twenty-nine point 9 two hundred sixty-two point 9 9 twenty-six point 2 9 9

OpenStudy (jojo4eva):

.32/8=0.04

OpenStudy (jojo4eva):

so the answer is 0.04

OpenStudy (anonymous):

ok can u answer 2

OpenStudy (anonymous):

omfg 0.04 i meant HAHAHAH this is why i never get full marks for easy papers

OpenStudy (jojo4eva):

11.9*22.1=262.99

OpenStudy (jojo4eva):

2.6299

OpenStudy (anonymous):

Which of the following is greatest? thirteen twenty-fifths zero point four five forty-three one hundredths twenty-nine fiftieths

OpenStudy (jojo4eva):

so its 262.99

OpenStudy (jojo4eva):

the answer to the second one is 262.99

OpenStudy (anonymous):

ok wat about 3

OpenStudy (jojo4eva):

what is the third one

OpenStudy (anonymous):

i'm careless to the extreme of ridiculousness really sorry bout that T_T 13x0.04=0.52 0.45 43x0.001=0.043 29x0.02=0.58

OpenStudy (anonymous):

wat about Which of the following is greatest? thirteen twenty-fifths zero point four five forty-three one hundredths twenty-nine fiftieths

OpenStudy (jojo4eva):

29/50= 0.58=58%

OpenStudy (anonymous):

What is the value of the expression 3x + x – 10 when x = 4? 58 19 11 6

OpenStudy (jojo4eva):

6

OpenStudy (anonymous):

Identify the place value of the zero that is not needed in the number one hundred twenty point nine four five zero. ones thousands ten-thousandths tenths

OpenStudy (anonymous):

Divide eight and one-sixth divided by four and two-thirds. two and one-half four-sevenths two and two-thirds one and three-fourths

OpenStudy (anonymous):

@kelliegirl33

OpenStudy (anonymous):

I need help with the last one pls

OpenStudy (anonymous):

@kelliegirl33

OpenStudy (jojo4eva):

u got it

OpenStudy (jojo4eva):

two and two thirds